# GFT: Graph Feature Tuning for Efficient Point Cloud Analysis

### v1

@@ Abstract
Parameter-efficient fine-tuning (PEFT) significantly reduces computational and memory costs by updating only a small subset of the model’s parameters, enabling faster adaptation to new tasks with minimal loss in performance. Previous studies have introduced PEFTs tailored for point cloud data, as general approaches are suboptimal. To further reduce the number of trainable parameters, we propose a point-cloud-specific PEFT, termed Graph Features Tuning (GFT), which learns a dynamic graph from initial tokenized inputs of the transformer using a lightweight graph convolution network and passes these graph features to deeper layers via skip connections and efficient cross-attention modules. Extensive experiments on object classification and segmentation tasks show that GFT operates in the same domain, rivalling existing methods, while reducing the trainable parameters. Code is at https://github.com/manishdhakal/GFT.
@@ 1 Introduction
Earlier methods of point cloud modeling include PointNet [20], PoinNet++ [21], and DGCNN [30]. These methods utilize the global and local features from point clouds in $(x,y,z)$ coordinates space, yielding significant performance in classification and segmentation. However, there has been evidence that these point cloud representation learning methods may not be optimal due to random initialization of weights [5, 19, 37]. One approach to fixing these issues is to perform self-supervised pretraining and full fine-tuning (FFT) for some downstream tasks. This has been demonstrated to be effective in various works [16, 17, 35].
@@
FFT of the pretrained models is expensive, as a separate copy of the fine-tuned model is required to be stored for every new additional tasks. It demands more storage expense for horizontal scaling of the tasks and reduces the degree of parameters sharing. To overcome this issue, various parameter-efficient fine-tuning (PEFT) methods—such as Adapters [11], and VPT [14],—have been developed for pretrained transformers [28] networks. These PEFTs have shown significant improvement in fine-tuning language and computer vision tasks. However, Zhou et al. [44] exhibited the sub-optimality of these generic PEFTs to adapt downstream tasks for point cloud analysis, implying a need for PEFTs specific to point cloud [38, 44]. IDPT [38] and DAPT [44] rose as point-cloud-specific PEFTs to fine-tune downstream tasks.
Learning graph features within the tokens of transformers has been an intuitive approach for fine-tuning transformer [28] models. The motivations for incorporating graph features stems from the transformer architecture itself, where self-attention can be interpreted as a fully connected, dynamic graph with a complexity of $\mathcal{O}(n^{2})$ over $n$ tokens, and the attention maps serve as edge weights. For efficiency, we can make this dynamic graph update sub-quadratic or linear with $k\mathcal{O}(n)$ complexity ($k\ll n$) via KNN graphs. Additionally, Figure 1 demonstrates the effectiveness of fine-tuning graph features, as it rewrites the attention maps to capture discriminative local geometric structures—effectively learning where to focus. In contrast, the attention maps from the pretrained model appear random and scattered, lacking a clear sense of focus.
@@
IDPT [38] employs a heavy graph features learner only in the model’s final layer. However, this approach has a significant drawback: the model is relatively heavy and sensitive to parameter budget reduction (as shown in Figure 2). Our goal is to design a lightweight and point-cloud-specific PEFT that can learn graph features for fine-tuning while being parameter-budget-friendly.
Figure 3: Full fine-tuning (FFT) vs. IDPT [38] vs. GFT. (classification task) (a) The whole model is fine-tuned in an end-to-end manner, including the encoder and the task head. (b) IDPT freezes the encoders and trains a heavy graph feature extractor at the last layer. (c) With light blocks, GFT extracts graph features from the earliest tokens of transformers and selectively injects features into the encoder.
With this study, we propose Graph Features Tuning (GFT) that incorporates additional graph features from the transformer’s input tokens to the deeper layers of the model, as shown in Figure 3. GFT introduces three modules that adapt downstream tasks: (1) Task-specific prompt [14] module that prepends learnable continuous token to token embedding space for higher degree of freedom for the graph. (2) EdgeConv [30] module has multiple layers of graph convolutional network (GCN) that embed locality information. (3) Cross-attention interaction module injects the graph features into the encoder. All of these modules are extremely small, updating merely 0.73 million parameters to fine-tune a pretrained transformer for classification.
@@ 2.1 Representation Learning of Point Clouds
Earlier self-supervised models used convolution-like operators [30, 32] to pass messages between points and their neighbours [40, 29, 41] for representation learning. With the rise of Transformers [28], Point Transformers [42] have been devised to encode a group of points as a single token, and multiple sequential tokens were used as inputs for stacked self-attention networks [19, 39, 5, 37, 22]. Following the principles of masked language modeling, most of the self-supervised point cloud models learn from masked point modeling that randomly masks points from the cloud and enforces the models to predict the missing signal. Point-BERT [37] learns to generate discrete point tokens containing meaningful information about the missing signals by using discrete Variational AutoEncoder (dVAE). Point-MAE [19], PointM2AE [39], and ReCon [22] have shown significant improvement in the pertaining process by reconstructing the missing information. ACT [5] learns the representation with self-supervision by employing cross-modal transfer of weights from a 2D image or language encoder. In our work, we use the pretrained models with architectures based on point transformers trained with different reconstruction methods and a variety of pertaining recipes.
@@ 2.2 Parameter Efficient Fine-Tuning for Point-Clouds
Parameter-efficient fine-tunings (PEFTs) modulate the representation of the pretrained models with addition of fewer learnable parameters for the downstream tasks. Adapters update pretrained models by training smaller blocks that runs in parallel to pretrained blocks [11, 25], effectively fine-tuning vision and language tasks [11, 33, 4, 12]. Prompt tuning, on the other hand, introduces learnable continuous tokens into the transformer’s embedding space while keeping the pretrained model frozen [14, 43, 26, 6, 1]. Although these PEFT methods have demonstrated success in vision and language domains, they fall short in point cloud analysis [44, 38]. To address this, two PEFT approaches have been specifically designed for point clouds: IDPT [38] and DAPT [44]. IDPT learns prompts from the last layer by generating a graph between tokens, while DAPT fine-tunes adapters by generating a dynamic scale for each token. In this work, we propose a PEFT that further optimizes efficiency for point cloud tasks.
@@ 2.3 Message Passing in Point Cloud
Message passing from neighboring data to generate local features is straightforward for text and 2D images because they inherently possess structural information. In these domains, convolutional operators can efficiently aggregate information from neighboring elements within the feed-forward network [9, 3, 15]. However, point clouds lack such structural information, as they are represented as an unordered set of xyzxyz coordinates. To enable convolution-like message passing, a graph-based structure is required to define relationships between points, typically using K-nearest neighbors (KNN) [21, 30]. In this work, we utilize EdgeConv [30] to encode locality features from the token space of the transformer, allowing for effective locality feature extraction.
@@ 3.1 Preliminary of Point Transformer
Point Transformer [42] encodes 3D points $x\in\mathbb{R}^{N\times 3}$ from $(x,y,z)$ space to $D$-dimensional latent space. It contains two major components: a point tokenizer that divides the entire $N$ points into $L$ tokens and $F$ (usually $F=12$) transformer encoder layers with self-attention mechanisms.
@@ 3.2.1 Task-specific Learnable Prompt
Learnable prompt tuning is one of the efficient fine-tuning methods used in the 2D-vision and language domains to learn task-specific features[14, 6]. This extremely parameter-efficient method injects learnable continuous prompts into the token embedding space of the transformer while the entire encoder remains frozen. Similarly, we inject these learnable prompts to the initial embeddings $E_{0}$ of Equation 1 between $T_{cls}$ and other tokens as:
@@ 3.2.2 EdgeConv Feature Extraction
EdgeConv [30] extracts features from the initial embedding matrix $E_{0}$ (excluding the $T_{cls}$ token) of the transformer pipeline. It is a component of multi-layered graph convolution functions for message passing that encodes features from the immediate neighbours. However, $E_{0}$ does not contain the intrinsic neighbourhood information of a token. Thus, we define the locality information of $i^{th}$ token $T_{i}\in\mathbb{R}^{1\times D}$ by grouping its K-nearest neighbours $\mathcal{N}(T_{i})\in\mathbb{R}^{K\times D}$. Distance from the neighbours to $T_{i}$ is attributed as edge features for KNN graphs, as shown in Figure 4 (b):
@@ 3.2.3 Cross-attention Interactions
Features from EdgeConv interact with the pretrained Point-Transformer to update the transformer’s features. Each interaction includes a cross-attention module [28] to fine-tune the pretrained features. To make the interaction lightweight, we implement sparse interaction to self-attention layers of point transformer. For instance, our implementation of $F=12$ layered encoders is subjected to $4$ interactions at $i\in\{1^{st},4^{th},7^{th},10^{th}\}$ self-attention layers.
@@
PointNet [20] CVPR’17 3.5 0.5 73.3 79.2 68.0 89.2
PointNet++ [21] NeurIPS’17 1.5 1.7 82.3 84.3 77.9 90.7
DGCNN [30] TOG’19 1.8 2.4 82.8 86.2 78.1 92.9
MVTN [8] ICCV’21 11.2 43.7 - - 82.8 93.8
PointNeXt [23] NeurIPS’22 1.4 1.6 - - 87.7 94.0
PointMLP [18] ICLR’22 13.2 31.4 - - 85.4 94.5
RepSurf-U [24] CVPR’22 1.5 0.8 - - 84.3 94.4
ADS [10] ICCV’23 - - - - 87.5 95.1
@@
OcCo [29] ICCV’21 22.1 4.8 84.85 85.54 78.79 92.1
MaskPoint [17] ECCV’22 22.1 - 89.70 89.30 84.60 93.8
Point-M2AE [39] NeurIPS’22 15.3 3.6 91.22 88.81 86.43 94.0
ReCon[22] ICML’23 43.6 5.3 94.15 93.12 89.73 93.9
Point-MAE [19] ECCV’22 22.1 4.8 90.02 88.29 85.18 93.8
Point-BERT [37] CVPR’22 22.1 4.8 87.43 88.12 83.07 93.2
ACT [5] ICLR’23 22.1 4.8 93.29 91.91 88.21 93.7
@@
IDPT [38] ICCV’23 1.70 (7.69%) 7.2 92.43 $\pm$ 0.27 91.43 $\pm$ 0.28 88.27 $\pm$ 0.23 93.00 $\pm$ 0.18
DAPT [44] CVPR’24 1.10 (4.97%) 5.0 91.81 $\pm$ 0.36 91.36 $\pm$ 0.76 87.56 $\pm$ 0.64 92.88 $\pm$ 0.21
@@
IDPT [38] ICCV’23 1.70 (7.69%) 7.2 91.98 $\pm$ 0.76 91.67 $\pm$ 0.78 86.91 $\pm$ 0.78 92.67 $\pm$ 0.17
DAPT [44] CVPR’24 1.10 (4.97%) 5.0 93.05 $\pm$ 0.59 91.67 $\pm$ 0.50 88.63 $\pm$ 0.19 92.49 $\pm$ 0.08
@@
IDPT [38] ICCV’23 1.70 (7.69%) 7.2 91.60 $\pm$ 0.39 91.60 $\pm$ 0.37 87.15 $\pm$ 0.39 93.07 $\pm$ 0.20
DAPT [44] CVPR’24 1.10 (4.97%) 5.0 91.15 $\pm$ 0.20 91.15 $\pm$ 0.52 86.97 $\pm$ 0.45 92.63 $\pm$ 0.21
@@
Table 1: Overall accuracy (OA% $\uparrow$) of classification on three variants of the ScanObjectNN [27] and the ModelNet40 [34]. #T-params and FLOPS (G) indicate the number of trainable parameters (in millions) and inference floating point operations, respectively. ScanObjectNN used the strong data augmentation mentioned by ACT [5].
@@ 4.1 Baseline PEFTs
To cover a wider spectrum of pretraining strategies for Point Transformer, we include three backbones: Point-BERT [37], Point-MAE [19], and ACT [5]. These foundation models were pretrained with ShapeNetCore from ShapeNet [2], containing $\sim$51K CAD models of 55 common object categories. We include the results from two point-cloud-specific PEFTs, IDPT [38] and DAPT [44], to compare with our method.

### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) significantly reduces computational and memory costs by updating only a small subset of the model’s parameters, enabling faster adaptation to new tasks with minimal loss in performance. Previous studies have introduced PEFTs tailored for point cloud data, as general approaches are suboptimal. To further reduce the number of trainable parameters, we propose a point-cloud-specific PEFT, termed Graph Features Tuning (GFT), which learns a dynamic graph from initial tokenized inputs of the transformer using a lightweight graph convolution network and passes these graph features to deeper layers via skip connections and efficient cross-attention modules. Extensive experiments on object classification and segmentation tasks show that GFT operates in the same domain, rivalling existing methods, while reducing the trainable parameters. Code is available at https://github.com/manishdhakal/GFT.
@@ 1 Introduction
Earlier methods of point cloud modeling include PointNet [qi2017pointnet], PoinNet++ [qi2017pointnet++], and DGCNN [wang2019dynamic]. These methods utilize the global and local features from point clouds in $(x,y,z)$ coordinates space, yielding significant performance in classification and segmentation. However, there has been evidence that these point cloud representation learning methods may not be optimal due to random initialization of weights [dong2023autoencoders, pang2022masked, yu2022point]. One approach to fixing these issues is to perform self-supervised pretraining and full fine-tuning (FFT) for some downstream tasks. This has been demonstrated to be effective in various works [liang2025pointmamba, liu2022masked, xie2020pointcontrast].
@@
FFT of the pretrained models is expensive, as a separate copy of the fine-tuned model is required to be stored for every new additional tasks. It demands more storage expense for horizontal scaling of the tasks and reduces the degree of parameters sharing. To overcome this issue, various parameter-efficient fine-tuning (PEFT) methods—such as Adapters [houlsby2019parameter], and VPT [jia2022visual],—have been developed for pretrained transformers [vaswani2017attention] networks. These PEFTs have shown significant improvement in fine-tuning language and computer vision tasks. However, zhou2024dynamic exhibited the sub-optimality of these generic PEFTs to adapt downstream tasks for point cloud analysis, implying a need for PEFTs specific to point cloud [zha2023instance, zhou2024dynamic]. IDPT [zha2023instance] and DAPT [zhou2024dynamic] rose as point-cloud-specific PEFTs to fine-tune downstream tasks.
Learning graph features within the tokens of transformers has been an intuitive approach for fine-tuning transformer [vaswani2017attention] models. The motivations for incorporating graph features stems from the transformer architecture itself, where self-attention can be interpreted as a fully connected, dynamic graph with a complexity of $\mathcal{O}(n^{2})$ over $n$ tokens, and the attention maps serve as edge weights. For efficiency, we can make this dynamic graph update sub-quadratic or linear with $k\mathcal{O}(n)$ complexity ($k\ll n$) via KNN graphs. Additionally, Figure 1 demonstrates the effectiveness of fine-tuning graph features, as it rewrites the attention maps to capture discriminative local geometric structures—effectively learning where to focus. In contrast, the attention maps from the pretrained model appear random and scattered, lacking a clear sense of focus.
@@
IDPT [zha2023instance] employs a heavy graph features learner only in the model’s final layer. However, this approach has a significant drawback: the model is relatively heavy and sensitive to parameter budget reduction (as shown in Figure 2). Our goal is to design a lightweight and point-cloud-specific PEFT that can learn graph features for fine-tuning while being parameter-budget-friendly.
Figure 3: Full fine-tuning (FFT) vs. IDPT [zha2023instance] vs. GFT. (classification task) (a) The whole model is fine-tuned in an end-to-end manner, including the encoder and the task head. (b) IDPT freezes the encoders and trains a heavy graph feature extractor at the last layer. (c) With light blocks, GFT extracts graph features from the earliest tokens of transformers and selectively injects features into the encoder.
With this study, we propose Graph Features Tuning (GFT) that incorporates additional graph features from the transformer’s input tokens to the deeper layers of the model, as shown in Figure 3. GFT introduces three modules that adapt downstream tasks: (1) Task-specific prompt [jia2022visual] module that prepends learnable continuous token to token embedding space for higher degree of freedom for the graph. (2) EdgeConv [wang2019dynamic] module has multiple layers of graph convolutional network (GCN) that embed locality information. (3) Cross-attention interaction module injects the graph features into the encoder. All of these modules are extremely small, updating merely 0.73 million parameters to fine-tune a pretrained transformer for classification.
@@ 2.1 Representation Learning of Point Clouds
Earlier self-supervised models used convolution-like operators [wang2019dynamic, wu2019pointconv] to pass messages between points and their neighbours [zhang2022pointdae, wang2021unsupervised, zhang2021self] for representation learning. With the rise of Transformers [vaswani2017attention], Point Transformers [zhao2021point] have been devised to encode a group of points as a single token, and multiple sequential tokens were used as inputs for stacked self-attention networks [pang2022masked, zhang2022point, dong2023autoencoders, yu2022point, qi2023contrast]. Following the principles of masked language modeling, most of the self-supervised point cloud models learn from masked point modeling that randomly masks points from the cloud and enforces the models to predict the missing signal. Point-BERT [yu2022point] learns to generate discrete point tokens containing meaningful information about the missing signals by using discrete Variational AutoEncoder (dVAE). Point-MAE [pang2022masked], PointM2AE [zhang2022point], and ReCon [qi2023contrast] have shown significant improvement in the pertaining process by reconstructing the missing information. ACT [dong2023autoencoders] learns the representation with self-supervision by employing cross-modal transfer of weights from a 2D image or language encoder. In our work, we use the pretrained models with architectures based on point transformers trained with different reconstruction methods and a variety of pertaining recipes.
@@ 2.2 Parameter Efficient Fine-Tuning for Point-Clouds
Parameter-efficient fine-tunings (PEFTs) modulate the representation of the pretrained models with addition of fewer learnable parameters for the downstream tasks. Adapters update pretrained models by training smaller blocks that runs in parallel to pretrained blocks [houlsby2019parameter, rebuffi2017learning], effectively fine-tuning vision and language tasks [houlsby2019parameter, wu2025slora, dhakal2024vlsm, hu2022lora]. Prompt tuning, on the other hand, introduces learnable continuous tokens into the transformer’s embedding space while keeping the pretrained model frozen [jia2022visual, zhou2022learning, shin2020autoprompt, gao2021making, adhikari2024tunevlseg]. Although these PEFT methods have demonstrated success in vision and language domains, they fall short in point cloud analysis [zhou2024dynamic, zha2023instance]. To address this, two PEFT approaches have been specifically designed for point clouds: IDPT [zha2023instance] and DAPT [zhou2024dynamic]. IDPT learns prompts from the last layer by generating a graph between tokens, while DAPT fine-tunes adapters by generating a dynamic scale for each token. In this work, we propose a PEFT that further optimizes efficiency for point cloud tasks.
@@ 2.3 Message Passing in Point Cloud
Message passing from neighboring data to generate local features is straightforward for text and 2D images because they inherently possess structural information. In these domains, convolutional operators can efficiently aggregate information from neighboring elements within the feed-forward network [he2016deep, conneau2016very, kalchbrenner2014convolutional]. However, point clouds lack such structural information, as they are represented as an unordered set of xyzxyz coordinates. To enable convolution-like message passing, a graph-based structure is required to define relationships between points, typically using K-nearest neighbors (KNN) [qi2017pointnet++, wang2019dynamic]. In this work, we utilize EdgeConv [wang2019dynamic] to encode locality features from the token space of the transformer, allowing for effective locality feature extraction.
@@ 3.1 Preliminary of Point Transformer
Point Transformer [zhao2021point] encodes 3D points $x\in\mathbb{R}^{N\times 3}$ from $(x,y,z)$ space to $D$-dimensional latent space. It contains two major components: a point tokenizer that divides the entire $N$ points into $L$ tokens and $F$ (usually $F=12$) transformer encoder layers with self-attention mechanisms.
@@ 3.2.1 Task-specific Learnable Prompt
Learnable prompt tuning is one of the efficient fine-tuning methods used in the 2D-vision and language domains to learn task-specific features[jia2022visual, gao2021making]. This extremely parameter-efficient method injects learnable continuous prompts into the token embedding space of the transformer while the entire encoder remains frozen. Similarly, we inject these learnable prompts to the initial embeddings $E_{0}$ of Equation 1 between $T_{cls}$ and other tokens as:
@@ 3.2.2 EdgeConv Feature Extraction
EdgeConv [wang2019dynamic] extracts features from the initial embedding matrix $E_{0}$ (excluding the $T_{cls}$ token) of the transformer pipeline. It is a component of multi-layered graph convolution functions for message passing that encodes features from the immediate neighbours. However, $E_{0}$ does not contain the intrinsic neighbourhood information of a token. Thus, we define the locality information of $i^{th}$ token $T_{i}\in\mathbb{R}^{1\times D}$ by grouping its K-nearest neighbours $\mathcal{N}(T_{i})\in\mathbb{R}^{K\times D}$. Distance from the neighbours to $T_{i}$ is attributed as edge features for KNN graphs, as shown in Figure 4 (b):
@@ 3.2.3 Cross-attention Interactions
Features from EdgeConv interact with the pretrained Point-Transformer to update the transformer’s features. Each interaction includes a cross-attention module [vaswani2017attention] to fine-tune the pretrained features. To make the interaction lightweight, we implement sparse interaction to self-attention layers of point transformer. For instance, our implementation of $F=12$ layered encoders is subjected to $4$ interactions at $i\in\{1^{st},4^{th},7^{th},10^{th}\}$ self-attention layers.
@@
PointNet [qi2017pointnet] CVPR’17 3.5 0.5 73.3 79.2 68.0 89.2
PointNet++ [qi2017pointnet++] NeurIPS’17 1.5 1.7 82.3 84.3 77.9 90.7
DGCNN [wang2019dynamic] TOG’19 1.8 2.4 82.8 86.2 78.1 92.9
MVTN [hamdi2021mvtn] ICCV’21 11.2 43.7 - - 82.8 93.8
PointNeXt [qian2022pointnext] NeurIPS’22 1.4 1.6 - - 87.7 94.0
PointMLP [ma2022rethinking] ICLR’22 13.2 31.4 - - 85.4 94.5
RepSurf-U [ran2022surface] CVPR’22 1.5 0.8 - - 84.3 94.4
ADS [hong2023attention] ICCV’23 - - - - 87.5 95.1
@@
OcCo [wang2021unsupervised] ICCV’21 22.1 4.8 84.85 85.54 78.79 92.1
MaskPoint [liu2022masked] ECCV’22 22.1 - 89.70 89.30 84.60 93.8
Point-M2AE [zhang2022point] NeurIPS’22 15.3 3.6 91.22 88.81 86.43 94.0
ReCon[qi2023contrast] ICML’23 43.6 5.3 94.15 93.12 89.73 93.9
Point-MAE [pang2022masked] ECCV’22 22.1 4.8 90.02 88.29 85.18 93.8
Point-BERT [yu2022point] CVPR’22 22.1 4.8 87.43 88.12 83.07 93.2
ACT [dong2023autoencoders] ICLR’23 22.1 4.8 93.29 91.91 88.21 93.7
@@
IDPT [zha2023instance] ICCV’23 1.70 (7.69%) 7.2 92.43 $\pm$ 0.27 91.43 $\pm$ 0.28 88.27 $\pm$ 0.23 93.00 $\pm$ 0.18
DAPT [zhou2024dynamic] CVPR’24 1.10 (4.97%) 5.0 91.81 $\pm$ 0.36 91.36 $\pm$ 0.76 87.56 $\pm$ 0.64 92.88 $\pm$ 0.21
@@
IDPT [zha2023instance] ICCV’23 1.70 (7.69%) 7.2 91.98 $\pm$ 0.76 91.67 $\pm$ 0.78 86.91 $\pm$ 0.78 92.67 $\pm$ 0.17
DAPT [zhou2024dynamic] CVPR’24 1.10 (4.97%) 5.0 93.05 $\pm$ 0.59 91.67 $\pm$ 0.50 88.63 $\pm$ 0.19 92.49 $\pm$ 0.08
@@
IDPT [zha2023instance] ICCV’23 1.70 (7.69%) 7.2 91.60 $\pm$ 0.39 91.60 $\pm$ 0.37 87.15 $\pm$ 0.39 93.07 $\pm$ 0.20
DAPT [zhou2024dynamic] CVPR’24 1.10 (4.97%) 5.0 91.15 $\pm$ 0.20 91.15 $\pm$ 0.52 86.97 $\pm$ 0.45 92.63 $\pm$ 0.21
@@
Table 1: Overall accuracy (OA% $\uparrow$) of classification on three variants of the ScanObjectNN [uy2019revisiting] and the ModelNet40 [wu20153d]. #T-params and FLOPS (G) indicate the number of trainable parameters (in millions) and inference floating point operations, respectively. ScanObjectNN used the strong data augmentation mentioned by ACT [dong2023autoencoders].
@@ 4.1 Baseline PEFTs
To cover a wider spectrum of pretraining strategies for Point Transformer, we include three backbones: Point-BERT [yu2022point], Point-MAE [pang2022masked], and ACT [dong2023autoencoders]. These foundation models were pretrained with ShapeNetCore from ShapeNet [chang2015shapenet], containing $\sim$51K CAD models of 55 common object categories. We include the results from two point-cloud-specific PEFTs, IDPT [zha2023instance] and DAPT [zhou2024dynamic], to compare with our method.